Acme Delivery is considering a proposal for new package tracking technology. The system has an estimated initial cost of $1.9 million and will require upgrades and maintenance of $140,000 each year. Acme estimates that improved tracking will save approximately $680,000 per year. Acme has a MARR of 15% per year, and the study period for this technology is 6 years, after which time Acme expects the entire system will need to be replaced Determine how sensitive the decision to invest in the system is to the estimates of initial investment cost and annual savings What is the breakeven life of the system? a. b. To FindF To Find Given P To Find F Given A F/A To Find P Given A P/A To Find A Given F A/F To Find A Given F Given P A/P 1.1500 06151 0.4380 0.3503 0.2983 F/P 1.1500 0.8696 086% 16257 10000 1.0000 0.4651 0.2880 02003 3225 1.7490 2.3131 3.0590 0456 21500 3.4725 .9934 1.5209 0.7561 0.6575 0.5718 04972 2 2832 2.8550 0.3759 351790.3269 0.2843 8.7537 11.0668 13.7268 0.2642 0.2404 02229 020% 0.1993 4.1604 0.0904 16.78584873 47716 0.0729 005% 20.3037 5.0188 0.0493
Answer #1

This is a capital budgeting question

Here cost on implementing new tracking technology are cash outoutflows while savings from implementing technology are cash inflows.

In year 0 cost on technology is initial cost of $1.9 million or $1,900,000.

For next six year cost is maintenance cost of $140,000

Therefore total cost on implementing this technology is $1,900,000 + 6 × 140,000

MARR or minimum average rate of return for ACME is 15%

Hence to calculate present value of costs we need to discount them at 15 %

Which is = (1,900,000)/ (1+15%)^0 + 140,000/(1+15%)^1 + 140,000/(1+15%)^2 +140,000/(1+15%)^3 +140,000/(1+15%)^4 +140,000/(1+15%)^5 +140,000/(1+15%)^6

As in table discount factor values are provided we will use same. Like for 1/1.15 value is .8696 for 1/1.15^2 value given is .7561. Hence we will use till 6 years as 6 years is total life of technology. Initial cost is done in 0 year or present year only so it will not be discounted.

So present value of total cost = 1,900,000 + 140,000 (0.8696 + 0.7561 + 0.6575 + 0.5718+ 0.4972 + 0.4323)

= 1,900,000 + 140,000× 3.7845 = 1,900,000 + 529,830 = 2,429,830.

Total saving per year from this technology is 680,000 per year.

As we have already calculated discount rates for cost same will apply to savings.

So total savings will be 680,000 × 3.7845= 2,573,460.

So net present value is 2,573,460 - 2,429,830 = 143,630 which is positive.

Hence at current initial investment and annual savings it is beneficial for ACME to implement new tracking technology.

A) If initial cost will be more than 1,900,000 + 143,630 = 2,043,630 then cost will be more than savings and it will not be beneficial to implement new technology.

Same way present value of annual savings should be minimum 2,429,830/ discount factor or 2,429,830/3.7845 = $642,048

Hence if annual savings would be below $642,048 at current initial investment and maintenance cost it will not be beneficial to implement new technology.

B) It we calculate total cost without discounting it will be 1,900,000 + 140,000× 6 = 2,740,000.

At annual savings of $680,000 breakeven period will b 2,740,000/680,000 = 4.029.

Hence, ACME will breakeven in 5 years after implementation of technology.

Thanks

Or annual savings should be minimum 2,429,830/